Tohban Report 2013-10-09

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 02 Oct 2013
End Date: 09 Oct 2013
Tohban: Hazel Bain
Tohban email:
Next Tohban: Nicole Duncan
List all reports


Solar Activity

Solar activity was low to moderate this week. There are currently 10 ARs on the disk, the most active of which is NOAA 11865 which produced an M2 flare in the last 24 hours. This region is in the center of the disk and looks like it will continue to produce a number of C class flares in the next day or so. STEREO B shows one further AR that will rotate onto the disk in the next few days.

How many GOES flares occurred?

 Flares above B, C, M, X class were     14    11     1     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were     11     8     1     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      5     2     0     0

There were RHESSI flares/GOES flares 148 / 26 over the time range 02-Oct-13 09-Oct-13

Memory Management

The SSR is emptying by the end of the daily pass set.

Spacecraft Status

The cold plate 1 and 2 are at about 116 K and 115 K respectively.

Data Gaps

No data gaps this week.

Detector issues

The recurrent noise issue in detector 9 is still present. There was one small spike in detector 9 at 2-oct-2013 18:15.

Other notes

Spacecraft Management

No operations this week

Decimation Normal/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Enabled, during entire eclipse
Night time data (rears) Enabled, during entire eclipse
Require extra passes? No?
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? None
Personal tools